Solve the system by substitution.
3x – y = –7
x + y = –9
Help!! I am pulling my hair out! :0) I need help with substitution, I am not understanding. Can someone please help me. Thanks.

Answer by stanbon(48568)

Solve the system by substitution.
3x – y = –7
x + y = –9
--------------
Solve either equation for y:
y=3x+7
Substitute that into the OTHER equation to get:
x+(3x+7)=-9
4x=-16
x=-4
---------
Substitute that into EITHER of the equations to solve for "x":
x+-4=-9
x=-5
----------
Answer: (-4,-5)
You could check this in the other quation:
3x – y = –7
3(-4)-(-5)=-7
12+5=7
7=7
---------
Cheers,
Stan H.